A charge of 1μC is situated at the origin of the xy plane. A 3μC charge is placed at (1,2), a -4μC charge is placed at (-3,-5) and a -5μC charge is placed at (2,-4). Determine the force felt by the 1μC due to the presence of the other charges.
1
Expert's answer
2021-09-13T11:01:58-0400

To analyze the system we have to identify the charges:


"q_1=+1\, \mu C=1\times10^{-6}\,C
\\ q_2=+3\, \mu C=3\times10^{-6}\,C
\\ q_3=-4\, \mu C=-4\times10^{-6}\,C
\\ q_4=-5\, \mu C=-5\times10^{-6}\,C"


We also proceed to define the direction and unit vectors:


"r_{1/2}=r_{1}-r_{2}=(0,0)-(1,2)
\\ r_{1/2}=(-1,-2)
\\ \widehat{r_{1/2}}=\dfrac{(-1,-2)}{\sqrt{1^2+2^2}}=(-\frac{1}{\sqrt{5}},-\frac{2}{\sqrt{5}})
\\ r_{1/3}=r_{1}-r_{3}=(0,0)-(-3,-5)
\\ r_{1/3} =(3,5)
\\ \widehat{r_{1/3}}=\dfrac{(3,5)}{\sqrt{3^2+5^2}}=(\frac{3}{\sqrt{34}},\frac{5}{\sqrt{34}})
\\ r_{1/4}=r_{1}-r_{3}=(0,0)-(2,-4)
\\ r_{1/4}=(-2,4)
\\ \widehat{r_{1/4}}=\dfrac{(-2,4)}{\sqrt{2^2+4^2}}=(-\frac{1}{\sqrt{5}},\frac{2}{\sqrt{5}})"


After that we proceed to use Coulomb's law ("F=k\cfrac{q_iq_j}{r^2_{ij}}\,\widehat{r_{ij}}" ) for each of the pairs (between charges q1 and q2, charges q1 and q3, and charges q1 and q4) and then we can find the total force felt by q1:


"F_T=F_{1/2}+F_{1/3}+F_{1/4}
\\ F_T=kq_1 \Big( \dfrac{q_2}{r^2_{1/2}} \, \widehat{r_{1/2}} +\dfrac{q_3}{r^2_{1/3}} \, \widehat{r_{1/3}}+\dfrac{q_4}{r^2_{1/4}} \, \widehat{r_{1/4}} \Big)
\\ F_T=(9\times10^9\frac{Nm^2}{C^2})(10^{-6}\,C) \Bigg( \dfrac{(3\times10^{-6}\,C)}{(1^2+2^2)\,m^2}(-\frac{1}{\sqrt{5}},-\frac{2}{\sqrt{5}}) +...
\\...+\dfrac{(-4\times10^{-6}\,C)}{(3^2+5^2)\,m^2} (\frac{3}{\sqrt{34}},\frac{5}{\sqrt{34}})+\dfrac{(-5\times10^{-6}\,C)}{(2^2+4^2)\,m^2}(-\frac{1}{\sqrt{5}},\frac{2}{\sqrt{5}}) \Bigg)"


We proceed with the substitution to confirm the vector that describes the force felt by q1:


"\\ F_T=(9\times10^{-3}\,N) \Big( (\frac{3}{5}) (-\frac{1}{\sqrt{5}},-\frac{2}{\sqrt{5}})+...
\\...+(-\frac{2}{17})(\frac{3}{\sqrt{34}},\frac{5}{\sqrt{34}})+(-\frac{1}{4})(-\frac{1}{\sqrt{5}},\frac{2}{\sqrt{5}}) \Big)
\\ F_T=(9\times10^{-3}\,N) \Big( (-\frac{3}{5\sqrt{5}},-\frac{6}{5\sqrt{5}})+...
\\...+(-\frac{6}{17\sqrt{34}},-\frac{10}{17\sqrt{34}})+(\frac{1}{4\sqrt{5}},-\frac{1}{2\sqrt{5}}) \Big)
\\F_T= (9\,mN) \Big( -\frac{3}{5\sqrt{5}}-\frac{6}{17\sqrt{34}}+\frac{1}{4\sqrt{5}}\,,...
\\...\,, -\frac{6}{5\sqrt{5}} -\frac{10}{17\sqrt{34}} -\frac{1}{2\sqrt{5}} \Big)
\\ F_T \approx (9\times10^{-3}\,N) \Big( -0.217,-0.861 \Big)
\\ F_T \approxeq \Big( -1.953,-7.750 \Big)\times10^{-3}\,N"


After we find the force vector, we proceed to calculate its magnitude:


"\\ F_T = \sqrt{ \Big( (-1.953\times10^{-3}N)^2+(-7.750\times10^{-3}N)^2) \Big)}
\\F_T=7.992\times10^{-3}\,N"


In conclusion, the force felt by the 1μC due to the presence of the other charges is approximately 7.992 X 10-3 N or 7.992 mN (miliNewtons).
Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.